A Note on Discriminability of Lambda Terms Svitan Gaborovič, Maribor, Slovenia

Keywords: Lambda calculus, lambda term, discriminator, separability of terms, discriminability, numeral AMS Subject classification: 03B40; 68Q20 Abstract: We prove that the orders of any discriminable set of lambda terms are bounded.